P of 4, A wire carrying I 1.5 A into the page is placed in the external magnetic field B 0.8 T, that makes an angle 58 with the ty direction (as shown). What is the magnitude and direction of the magnetic force F (with respect to the +x axis) on the wire, if the length of the wire segment immersed in the field is 0.7 m? 58 inntaine a 100 ohm resistor, a 10 puF capacitor, and a 500 mH inductor that are of the 100 V (peak). Find the reactance the source
Answer #1

Magnetic force on a current carrying wire in a magnetic field is given by

F = I(L*B) = ILBsin(\Theta ) = 1.5*0.7*0.8*sin(90^{0}) = 0.84 N

This force will be perpendicular to both the Magnetic field and the current direction and according to right hand thumb rule, the force will be directed at 580 with respect to +x axis.